OpenStudy (anonymous):

What is the range of the quadratic function f(x) = -5x2 + 30x + 2?

OpenStudy (mertsj):

Find the vertex. Since this parabola is concave downward, the range will be all y values less than or equal to the y coordinate of the vertex.

OpenStudy (anonymous):

(-oo,2]

OpenStudy (mertsj):

The vertex is (3,47)

OpenStudy (anonymous):

Vertex ( h, k) h = - b/ 2a = 30/10 = 3

OpenStudy (anonymous):

f ( 3) = 47 => Range ( -∞, 47)

OpenStudy (mertsj):

So the range is all real numbers less than or equal to 47

OpenStudy (anonymous):

(-3.5,47.08)